OpenStudy (anonymous):

what is the solution to these lines? 5x-3y=7 3x+5y=10

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (radar):

-3y=-5x+7 y=5/3 x-7/3 m=5/3 5y=-3x+10 y=-3/5+2 m=-3/5 Since the slope are a negative reciprocaL of each other, the two lines are perpendicular.

OpenStudy (radar):

The two formulas now in slope intercept form can provide you the rest of the information. As always verify this work.